# Leveraging Graph Neural Networks and Multi-Agent Reinforcement Learning for Inventory Control in Supply Chains

### v1

@@ 2.1 Single Agent Reinforcement Learning - PPO
Proximal Policy Optimisation (PPO) is a popular first-order, on-policy [^1] Technically, PPO still employs off-policy corrections, meaning it reuses samples collected during training, but it does not explicitly use a replay buffer. PPO updates the policy network using a surrogate objective function that constrains the policy update to be within a certain proximity of the previous policy. This avoids the need to store and sample experiences single agent reinforcement learning method. PPO is an actor-critic algorithm where a policy $\pi_{\theta}$ and value function $V_{\phi}(s)$ are two separate neural networks parameterised by $\theta$ and $\phi$ respectively. In actor critic algorithms like PPO, $V_{\phi}(s)$ is introduced to reduce the variance but may introduce a bias in $\pi_{\theta}$. The two variations of PPO are using a penalty function or a clipping function. The later is known to be crucial for its performance as the clipping function constraints the ratio between the new and old policy within a certain range to prevent large policy updates that may lead to instability.
@@
where $\pi_{\theta}(a_{t}|s_{t})$ represents the probability of taking actions $a_{t}$ in state $s_{t}$ under the current policy parameterized by $\theta$, $\pi_{\theta_{old}}(a_{t}|s_{t})$ is fixed during the policy update step and represents the probability of taking action $a_{t}$ in state $s_{t}$ under the old policy (from the previous iteration) parameterized by $\theta_{\text{old}}$, $a_{t}$ and $s_{t}$ are the action and state respectively taken at time step $t$, $\epsilon$ represents a hyperparameter that determines how much the new policy can deviate from the old policy and $\text{clip}(\cdot,1-\epsilon,1+\epsilon)$ is the clipping function that constraints the ratio of the new to old policy’s probabilities within a certain range which prevents the policy from making overly large updates that could lead to instability. Finally, $A_{t}$ is the advantage function defined as:
@@
where $R_{t}$ and $V_{\phi}(s_{t})$ is the reward and value function at time step $t$ respectively and $\gamma$ is the discount factor. The general advantage estimation (GAE), used to compute the advantage, is given by:
@@ 2.2 Multi-Agent Reinforcement Learning
Two multi-agent variants of the popular single agent PPO algorithm are possible: IPPO and MAPPO.
Independent PPO (IPPO) IPPO is an independent learning algorithm which breaks down a problem with $n$ agents into $n$ decentralised single agent problems. A value function, $V^{i}_{\phi}(s)$, and policy, $\pi^{i}_{\theta}$, are present for each agent in IPPO, taking local inputs. Despite showing good overall performance in certain multi-agent settings, IPPO can lead to non-stationarity in the environment. This occurs because each agent’s policy is updated simultaneously which affects the state transition probability, $p(s^{\prime},r|s,a^{i},\pi)$, which becomes non-stationary. Therefore, theoretically, the convergence of the Bellman Equation, shown in Equation 4 is not guaranteed.
$V^{\pi^{i}}(s)=\sum_{a}\pi^{i}(a^{i}|s)\sum_{s^{\prime},r}p(s^{\prime},r|s,a^{i},\pi)(r+v_{\pi^{i}}(s^{\prime}))$ (4)
@@
$\mathcal{L}(o^{i},s,a^{i};\mathbf{a^{-}},\theta_{k},\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\min\left(\frac{\pi_{\theta^{i}}(a^{i}|o^{i})}{\pi_{\theta_{\text{old}}^{i}}(a^{i}|o^{i})}A^{\pi_{\theta_{\text{old}}^{i}}}(o^{i},s,\mathbf{a^{-}}),\right.\right.$
@@
Where $a$ is the current agent action, $\mathbf{a^{-}}$ is the action set of all agents, $s$ is the global state, $o$ is the local observation, The advantage function $A^{\pi_{\theta_{\text{old}}^{i}}}$ is computed using the GAE method in a similar manner to Equation 3.
However, when GAE is applied in a multi-agent setting with a shared value function, the advantage estimated for each agent can be identical. In MAPPO, this is always true in fully cooperative cooperative environments where all agents share a common reward function and experience the same state transitions. This makes it challenging to accurately quantify and distinguish the unique contribution of each individual agent to the overall performance, even though the policy gradient considers the actions of all agents when updating the policy for each agent. This is known as implicit multi-agent credit. In contrast, in IPPO, where each agent has its own independent value function, the advantage estimates differ across agents, even in cooperative environments. This is because each agent calculates its own GAE based on its individual observations, rewards, and value function.
$\frac{\delta\mathcal{L}}{\delta\pi^{i}(a_{t}^{i}|s_{t})}\propto\mathbb{E}\left[A^{\pi_{\theta_{k}}}\left(o,s,\mathbf{a^{-}}\right)\right]$ (6)
An additional complication is that for a large number of agents, the number of possible joint actions becomes vast, so exploring the joint action space and create enough excitation for all agents to compute the true gradient is impractical. Therefore, learning algorithms rely on sampling techniques to estimate the gradient which may not adequately explore the joint agent space. This may lead to the problem of policy overfitting in cooperative multi-agent environments. Several studies have been conducted to reduce credit assignment issues through techniques like reward shaping (Zhou et al., 2020), employing individual critics (Hernandez-Leal et al., 2019) and communication protocols (Feng et al., 2022a). Other methods have been developed to tackle the effects of credit assignment by reducing policy overfitting with exploration bonuses (Yarahmadi, 2023) or regularized policy gradients (Liu et al., 2021).
@@
Figure 1: Schematic showcasing the Centralised Training Decentralised Execution framework
@@ 3 Methodology
In this section, we outline our approach for a graph-based multi-agent PPO algorithm. First, we discuss the integration of graph neural networks (GNNs) into our framework. Next, we illustrate how noise injection is utilized as a regularizer within the value function in our approach. Finally, we define the inventory management problem, outlining its key components and objectives.
3.1 P-GCN-MAPPO
In environments where the graph structure can be leveraged, Graph Neural Networks (GNNs) are commonly integrated into RL frameworks. GNNs were developed to efficiently leverage the structure and properties of graphs. GNNs operate on graph-structured data and are able to capture complex relationships and dependencies inherent in graphs. In this work, we use Graph Convolutional Networks (GCNs) combined with Multi-Agent Proximal Policy Optimization (MAPPO) and a Pooling strategy, hence P-GCN-MAPPO. In Section 4 we conduct computational experiments to analyze the different components of this methodology in an effort to distill their contribution to the overarching framework. GCNs update the representation of a node by aggregating and transforming the features of its neighbouring nodes and itself. This allows the model to capture and propagate local information, effectively learning patterns and dependencies from the graph structure of a supply chain. This ability to capture complex relationships is particularly beneficial for supply chain problems as understanding the intricate relationships between various entities can significantly optimize decision-making.
The adjacency matrix, $A\in\mathbb{R}^{N\times N}$, is a $N\times N$ matrix which is used to express the directed graph topology where $N$ is the number of nodes (or vertices) in the graph. In this matrix, $A_{ij}=1$ indicates that there is a connection between node $i$ and node $j$. Each node $i$ is associated with a node feature $x_{i}$ which encapsulates information specific to that node where $x_{i}\in\mathbb{R}^{D}$ where $D$ represents the dimensionality of the features vector. These individual node features collectively form the node feature matrix, $X\in\mathbb{R}^{N\times D}$. At each time step, the node feature matrix, $X$, captures the evolving state of the graph. Both $A$ and $X$ are fed into a graph convolution layer, allowing the model to capture relational information between nodes.
$f(X,A)\coloneqq\sigma(D^{-\frac{1}{2}}(A+I)D^{-\frac{1}{2}}XW)$ (7)
where $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix, $I\in\mathbb{R}^{N\times N}$ is the identity matrix, $D\in\mathbb{R}^{N\times N}$ is the degree matrix of $A+I$, $X\in\mathbb{R}^{N\times D}$ is the node feature matrix, $W\in\mathbb{R}^{D\times W^{\prime}}$ is the layer’s weights where $W^{\prime}$ is the number of output features and $\sigma(.)$ is the activation function (e.g. ReLU). This results in an embedded vector for each node, $h_{i}\in\mathbb{R}^{W^{\prime}}$. In this work, three convolution layers are used where the embedded vector at each node, $h_{i}$, is the input of the next layer. This is described mathematically as:
$H_{1}$ $\coloneqq f_{W_{1}}(X,A)$ (8)
$H_{2}$ $\coloneqq f_{W_{2}}(H_{1},A)$ (9)
$H_{3}$ $\coloneqq f_{W_{3}}(H_{2},A)$ (10)
where $H_{1}\in\mathbb{R}^{N\times W_{1}^{\prime}},H_{2}\in\mathbb{R}^{N\times W_{2}^{\prime}},H_{3}\in\mathbb{R}^{N\times W_{3}^{\prime}}$ are the embedded node matrices at layers 1,2 and 3 and $W_{1},W_{2},W_{3}$ are the weight matrices that parameterize each layer. The terms $W_{1}^{\prime},W_{2}^{\prime},W_{3}^{\prime}\in\mathbb{Z}^{+}$ are defined as positive integers representing the number of output features for each layer respectively. These values are hyperparameters that determine the dimensionality of the embedded vectors produced by each layer.
The resulting latent representations are then passed to our centralized value function $V_{\phi}(s)$, a fully connected network parameterized by $\phi$. This function assesses the value of actions, aiding in variance reduction of the policy $\pi_{\theta}$, another fully connected network parameterized by $\theta$, which maps states to actions. To train the policies, we employ Multi-Agent Proximal Policy Optimization (MAPPO). Mini-batches are sampled, and the objective is to minimize the loss function for each batch shown in Equation 2.2 which is used to update the parameters, $\theta$ and $\phi$.
However, as the number of agents increases, the dimensionality of the node feature changes. To address this, we propose integrating global mean pooling within our framework, ensuring dimensionality remains constant as the number of agents increases.
Pooling operators in graphs were inspired by pooling methods in Convolutional Neural Networks (CNNs). Therefore, instead of simply concatenating all the latent hidden representations and feeding them directly into our central value function, we enhance computational efficiency by employing the global mean pool operator. The global mean pooling operator aggregates information across all nodes in a graph by taking the average of the node features. The GNN outputs hidden feature matrix, $H\in\mathbb{R}^{N\times W_{3}^{\prime}}$ , can be denoted as $H=\{h_{1},h_{2},\dots,h_{N}\}$ where $N$ is the number of nodes and $h_{i}$ is the GNN output feature vector for node $i$. The global mean pooling operator computes the mean of each feature across all nodes.
Global Mean $\text{Global Mean Pool}(H)=\frac{1}{N}\sum_{i=1}^{N}h_{i}$ (11)
This effectively reduces the input dimensions to our critic from $N\times W_{3}^{\prime}$ to $1\times W_{3}^{\prime}$. This is shown in Figure 2. While this dimensionality reduction improves computational efficiency, there’s a potential drawback: it may result in the loss of specific spatial information. However, since each individual actor still leverages local information for each node, the loss is confined to the critic. Moreover, the critic might not require the full level of spatial detail to assess the overall value of a state. The high-level features extracted through the global mean pooling mechanism might suffice. This idea has been explored in several studies such as Fujimoto et al. (2018) who highlights the critic’s focus on good enough value approximation for policy improvement, suggesting it may not need full state representation and Lyu et al. (2023) which proved that centralized critics may not be beneficial; particularly state-based critics can introduced unexpected bias. This aligns with the idea that the critic might not need a perfect representation but an accurate enough estimate for policy guidance. Finally, the global mean pooling mechanisms can also help mitigate the risk of overfitting, a common problem in RL where the model performs well on training data but poorly on unseen data. By reducing the number of features, the model focuses on the most relevant aspects, potentially leading to better generalization.
3.2 Noisy P-GCN-MAPPO
The problem of policy over-fitting occurs as the estimated advantage is the same for all agents leading to the lack of credit assignment. Theoretically, this can be solved by ensuring the shared advantage value does not affect other agents by decomposing the centralized advantage value for each agent. However, given the nature of multi-agent systems, this is impractical. Therefore, noise can be introduced as a regularize to reduce bias in advantage values. Noise has been extensively studied and used in several forms across different components of reinforcement learning, including action space exploration and observation perturbation. In this work, we focus on introducing noise specifically in the value function which then propagates to the estimation of advantage values as shown in Figure 3. This approach offers several potential benefits and implications:
  1. 1.
  2. 2.
We sample a Gaussian noise,
$\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2})$ (12)
$V_{\phi}(s)\leftarrow V_{\phi}(s)+\epsilon_{n}$ (13)
@@ 3.3 Problem Statement
The cooperative nature of the MARL scenario that we consider means agents coordinate towards a common goal, receiving a shared reward. This sequential decision making problem is modelled as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP), which can be defined as a tuple $\left\langle\mathcal{S},\mathbf{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma\right\rangle.$ $\mathcal{S}$ is the set of all valid states, $\mathbf{A}\coloneqq\mathbf{A}^{1}\times\dots\times\mathbf{A}^{N}$ is the joint action space where $\mathbf{A}^{i}$ is the set of actions available for each agent $i$, $\Omega\coloneqq\Omega^{1}\times\dots\times\Omega^{N}$ is the joint observation space and $N$ denotes the number of agents. At each time step, each agent $i$ executes action ${a}^{i}\in\mathcal{A}^{i}\subseteq\mathbb{R}^{n_{a}}$ with a joint action $\textbf{a}=\left\langle a^{1},\dots,a^{N}\right\rangle$ and transitions from state $s\ \in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ to $s^{\prime}\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ with state transition probability $P(s^{\prime}|s,\textbf{a})=\mathcal{T}(s,\textbf{a},s^{\prime})$. Each agent $i$ receives observation $o^{i}\in\Omega^{i}\subseteq\mathbb{R}^{n_{o^{i}}}$ determined by $\mathcal{O}(s^{\prime},i)$ which maps the new state $s^{\prime}\in\mathcal{S}$ to an observation $o^{i}\in\Omega^{i}$ for each agent $i$. In other words, the observation function $\mathcal{O}:\mathcal{S}\times\{1,\dots,N\}\rightarrow\Omega$ provides each agent $i$ with a local observation $o^{i}$ based on the next state $s^{\prime,i}$. The joint observation can be defined as $\textbf{o}=\left\langle o^{1},\cdots,o^{N}\right\rangle$ and each agent shares the same reward function $\mathcal{R}(s,\textbf{a})\in\mathbb{R}$. Each agent has policy $\pi^{i}$ and the joint policy is denoted as $\mathbf{\pi}=\ \left\langle\pi^{1},\cdots,\ \pi^{N}\right\rangle$. The optimal joint policy is found through maximising the joint expected reward $\mathbb{E}\left[\sum_{t=0}^{t=T}\gamma^{t}r_{t}\right]$ where $r_{t}\ =\ \mathcal{R}\left(s_{t},\ a_{t}\right)\ \in\mathbb{R}$ at each time step $t$ where $T$ is the time horizon and the discount factor $\gamma\in\left[0,1\right]$.
@@
$b^{i}=\sum_{j\in\mathcal{D}_{i}}b^{i_{d}}$ Total backlog of node $i$ from downstream nodes
$g^{i}=\sum_{j\in\mathcal{D}_{i}}g^{i_{d}}$ Total shipment of node $i$ to downstream nodes
@@
Equation (LABEL:eqn:_optimisation) is the objective function that maximizes total profit across the supply chain system. The system is treated as a collaborative framework, where all agents in the network share a common objective function, reward $r_{t}$. Equation (LABEL:eqn:_inventory) and (LABEL:eqn:_backlog) show how the inventory and backlog are updated over time. Equations (LABEL:eqn:_sales_constraint_2) and (LABEL:eqn:_sales_constraint_1) restrict the quantity of goods a node can ship downstream, ensuring it does not exceed the on-hand inventory or the downstream demand and backlog. Equation (LABEL:eqn:_acquisition) and LABEL:eqn:_factory capture the lead time of a shipment, indicating that goods shipped to node $i$ will take $\tau^{i}$ periods to reach the downstream stage. The amount of inventory hold or ordered is also constrained to a maximum value through Equations LABEL:eqn:_limits. The interaction of the different flows between two nodes can also be seen in Figure 4.
@@
subject $\begin{split}&\text{subject to}\end{split}$
@@
The demand and lead time are uncertain, modelled as a Poisson random variable with parameter $\lambda$ and $\mu$ respectively. Therefore, the probability of $k$ demands in a time period $t$ is given by:
@@
Moreover, inventory management system can be modelled as a graph $G=(W,E)$ where $W$ represents different entities in the supply chain (e.g. factories, manufacturers, warehouses), $E$ denotes the relationship between these entities, and each node $i\in W$ corresponds to an agent responsible for managing inventory at that entity’s location. Each node $i$ has a node feature $x^{i}\in X$, representing the observation space $o^{i}$ for each agent. The observation set for each agent $i$ is $o^{i}=[v,b,p,d^{-1},\dots,d^{-M},o_{r}^{-1},\dots,o_{r}^{-M}]$ where $o^{i}\ \in\Omega^{i}$. In the observation set, a new variable is introduced, $p$, which is the pipeline inventory equal to the sum of order replenishment that has not yet arrived at the node from other upstream nodes. To mitigate the problem of partial observability, we include demand history and order history up to $M$ time-steps in the past where $M$ is a hyperparameter. While we chose not to incorporate Recurrent Neural Networks (RNNs) in this work, which could effectively handle sequential data and temporal dependencies, we acknowledge their potential as an alternative approach. Our decision to include a fixed window of past observations instead allows for a simpler implementation while still capturing relevant historical context.
@@
However, this inclusion of historical data introduces a violation of the Markov property, which states that the future state of the system only depends on the current state and not on the sequence of events that preceded it. However, real-world decision-making processes rarely exhibit perfect Markovian behavior. Therefore, including historical data is a well-established approach to augment the observation space in partially observable environments Liu et al. (2022a); Mousa et al. (2023); Uehara et al. (2022).
The neighbourhood of a node $i$, denoted by $M(i)$ is defined as the set of neighbouring nodes $j$ connected to node $i$ via edges in $E$. Mathematically, this can be represented as $M(i)=\{j|(i,j)\in E\}$. For each node $i$, the features of its neighbours $j$ are represented as $x_{j}^{i}$, capturing the relevant observations of the neighbouring agents.

### v2

@@ 2.2 Single Agent Reinforcement Learning - PPO
Proximal Policy Optimisation (PPO) is a popular first-order, on-policy [^1] Technically, PPO employs off-policy corrections, meaning it reuses samples collected during training, but it does not explicitly use a replay buffer. PPO updates the policy network using a surrogate objective function that constrains the policy update to be within a certain proximity of the previous policy. This avoids the need to store and sample experiences single agent reinforcement learning method. PPO is an actor-critic algorithm where a policy $\pi_{\theta}(a|s)$ and value function $V_{\phi}(s)$ are two separate neural networks parameterised by $\theta$ and $\phi$ respectively. In actor-critic algorithms like PPO, $V_{\phi}(s)$ is introduced to reduce the variance but may introduce a bias in $\pi_{\theta}$. The two variations of PPO are using a penalty function or a clipping function. The later is known to be crucial for its performance as the clipping function constraints the ratio between the new and old policy within a certain range to prevent large policy updates that may lead to instability.
@@
where $\pi_{\theta}(a_{t}|s_{t})$ represents the probability of taking actions $a_{t}$ at state $s_{t}$ under the current policy parameterized by $\theta$, $\pi_{\theta_{old}}(a_{t}|s_{t})$ is fixed during the policy update step and represents the probability of taking action $a_{t}$ at state $s_{t}$ under the old policy (from the previous iteration) parameterized by $\theta_{\text{old}}$, $a_{t}$ and $s_{t}$ are the action and state respectively taken at time step $t$, $\epsilon$ represents a hyperparameter that determines how much the new policy can deviate from the old policy and $\text{clip}(\cdot,1-\epsilon,1+\epsilon)$ is the clipping function that constraints the ratio of the new to old policy’s probabilities within a certain range which prevents the policy from making overly large updates that could lead to instability. Finally, $A_{t}$ is the advantage function defined as:
@@
where $R_{t}$ and $V_{\phi}(s_{t})$ are the reward and value function at time step $t$ respectively and $\gamma$ is the discount factor. The advantage function, $A_{t}$, quantifies how much better or worse the action taken at time $t$ is compared to the expected value of being in the state, effectively measuring the relative advantage of an action in improving future rewards. The general advantage estimation (GAE), used to compute the advantage, is given by:
@@ 2.3 Multi-Agent Reinforcement Learning
Two multi-agent variants of the popular single agent PPO algorithm are: IPPO and MAPPO.
Independent PPO (IPPO) IPPO is an independent learning algorithm which breaks down a problem with $n$ agents into $n$ decentralized single agent problems. A value function, $V^{i}_{\phi}(s)$, and policy, $\pi^{i}_{\theta}$, are present for each agent in IPPO, taking local inputs. Despite showing good overall performance in certain multi-agent settings, IPPO can lead to non-stationarity in the environment. This occurs because each agent’s policy is updated simultaneously which affects the state transition probability, $p(s^{\prime},r|s,a^{i},\pi)$, which becomes non-stationary. Therefore, the convergence of the Bellman Equation, shown in Equation 4 is not guaranteed and presents convergence problems in practice.
$V^{\pi^{i}}(s)=\sum_{a}\pi^{i}(a^{i}|s)\sum_{s^{\prime},r}p(s^{\prime},r|s,a^{i},\pi)(r+v_{\pi^{i}}(s^{\prime}))\quad\text{for }i=1,\ldots,n_{a}$ (4)
@@
$\mathcal{L}(o^{i},s,a^{i};\mathbf{a^{-}},\theta_{k},\theta)$ $=\frac{1}{n_{a}}\sum_{i=1}^{n_{a}}\mathbb{E}\left[\min\left(\frac{\pi_{\theta^{i}}(a^{i}|o^{i})}{\pi_{\theta_{\text{old}}^{i}}(a^{i}|o^{i})}A^{\pi_{\theta_{\text{old}}^{i}}}(o^{i},s,\mathbf{a^{-}}),\right.\right.$
@@
Where $a$ is the current agent action, $\mathbf{a^{-}}$ is the concatenated action of all agents, $s$ is the global state, $o$ is the local observation, The advantage function $A^{\pi_{\theta_{\text{old}}^{i}}}$ is computed using the GAE method in a similar manner to Equation 3.
However, when GAE is applied in a multi-agent setting with a shared value function, the advantage estimated for each agent can be identical. In MAPPO, this is always true in fully cooperative environments where all agents share a common reward function and experience the same state transitions. This makes it challenging to accurately quantify and distinguish the unique contribution of each individual agent to the overall performance, even though the policy gradient considers the actions of all agents when updating the policy for each agent. This is known as implicit multi-agent credit. In contrast, in IPPO, where each agent has its own independent value function, the advantage estimates differ across agents, even in cooperative environments. This is because each agent calculates its own GAE based on its individual observations, rewards, and value function.
$\frac{\delta\mathcal{L}}{\delta\pi^{i}(a_{t}^{i}|s_{t})}\propto\mathbb{E}\left[A^{\pi_{\theta_{old}}}\left(o,s,\mathbf{a^{-}}\right)\right]$ (6)
An additional complication is that for a large number of agents, the number of possible joint actions becomes vast, so exploring the joint action space and create enough excitation for all agents to compute the true gradient is impractical. Therefore, learning algorithms rely on sampling techniques to estimate the gradient which may not adequately explore the joint agent space. This may lead to the problem of policy overfitting in cooperative multi-agent environments. Several studies have been conducted to reduce credit assignment issues through techniques like reward shaping (Zhou et al., 2020b), employing individual critics (Hernandez-Leal et al., 2019) and communication protocols (Feng et al., 2022a). Other methods have been developed to tackle the effects of credit assignment by reducing policy overfitting with exploration bonuses (Yarahmadi, 2023) or regularized policy gradients (Liu et al., 2021).
@@
(a) Abstract Representation (Black-Box)
(b) Parametrized by Neural Networks
Figure 2: Schematic showcasing the Centralized Training Decentralized Execution framework, demonstrating how agents undergo collaborative training in a centralized manner (blue dashed line) while executing actions independently in their respective environments (red dashed line).
@@ 3 Methodology
In this section, we outline our approach for a graph-based multi-agent PPO algorithm. First, we define the inventory management problem, outlining its key components and objectives. Next, we discuss the graph representation of our problem. Finally, we describe our proposed framework with the integration of graph neural networks (GNNs) and illustrate how noise injection is utilized as a regularizer within the value function in our approach.
@@ 3.1 Problem Statement
$b^{i}=\sum_{j\in\mathcal{D}_{i}}b^{i_{d}}$ Total backlog of node $i$ from downstream nodes $j$
$g^{i}=\sum_{j\in\mathcal{D}_{i}}g^{i_{d}}$ Total shipment of node $i$ to downstream nodes $j$
@@
Equation (LABEL:eqn:_optimisation) is the objective function that maximizes total profit across the supply chain system. The system is treated as a collaborative framework, where all agents in the network share a common objective function, reward $\mathcal{R}_{t}$. The decision variables in this optimization include $o_{r}^{i}[t]$ which represents the order quantity at node $i$ at time $t$. Other decision-related quantities, like inventory and backlog, are derived from this order quantity based on the system’s dynamics, as captured by the relevant constraints. Equation (LABEL:eqn:_inventory) and (LABEL:eqn:_backlog) show how the inventory and backlog are updated over time. Equations (LABEL:eqn:_sales_constraint_2) and (LABEL:eqn:_sales_constraint_1) restrict the quantity of goods a node can ship downstream, ensuring it does not exceed the on-hand inventory or the downstream demand and backlog. Equation (LABEL:eqn:_acquisition) and (LABEL:eqn:_factory) capture the lead time of a shipment, indicating that goods shipped to node $i$ will take $\tau^{i}$ periods to reach the downstream stage. The amount of inventory hold or ordered is also constrained to a maximum value through Equations (LABEL:eqn:_limits). The interaction of the different flows between two nodes can also be seen in Figure 3.
The overarching goal is to maximize the net profit generated across all nodes and all time periods. We can formalize the optimization problem as follows:
@@
subject $\begin{split}&\text{subject to:}\end{split}$
@@
In our model, we take into account exogenous forms of uncertainty on the demand and lead time , modeled as a Poisson random variable with parameter $\lambda_{\text{d}}$ and $\lambda_{\text{l}}$ respectively. While inventory and backlog are treated deterministically to simplify the analysis and focus on long-term system performance, the framework indirectly incorporates the effects of uncertainty by modeling demand and lead time as stochastic variables - two key sources of uncertainty. This approach strikes a balance between computational efficiency and the need to capture essential uncertainties. Therefore, the probability of $k$ demands in a time period $t$ is given by:
@@
Moreover, we model the inventory control problem as a sequential decision-making problem where each node is modeled as a separate agent. The goal is to determine the optimal order quantity for each supply chain entity at each time step, ensuring coordination across agents while maximizing overall profit. In the context of multi-agent reinforcement learning, we model the supply chain as cooperative which means agents coordinate towards a common goal, receiving a shared reward. This sequential decision making problem is modelled as an extension of the MDP framework introduced in Section 2, adapted to a multi-agent and partially observable setting as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP), which can be defined as a tuple $\left\langle\mathcal{S},\mathbf{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma\right\rangle.$ $\mathcal{S}$ is the set of all valid states representing the joint state space shared by all agents, $\mathbf{A}\coloneqq\mathbf{A}^{1}\times\dots\times\mathbf{A}^{n_{a}}$ is the joint action space where $\mathbf{A}^{i}$ is the set of actions available for each agent $i$, $\Omega\coloneqq\Omega^{1}\times\dots\times\Omega^{n_{a}}$ is the joint observation space and $n_{a}$ denotes the number of agents. At each time step, each agent $i$ executes action ${a}^{i}\in\mathcal{A}^{i}\subseteq\mathbb{R}^{n_{a^{\prime}}}$ with a joint action $\textbf{a}=\left\langle a^{1},\dots,a^{n_{a}}\right\rangle$ and transitions from state $s\ \in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ to $s^{\prime}\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ with state transition probability $P(s^{\prime}|s,\textbf{a})=\mathcal{T}(s,\textbf{a},s^{\prime})$. Each agent $i$ receives observation $o^{i}\in\Omega^{i}\subseteq\mathbb{R}^{n_{o}}$ determined by $\mathcal{O}(s^{\prime},i)$ which maps the new state $s^{\prime}\in\mathcal{S}$ to an observation $o^{i}\in\Omega^{i}$ for each agent $i$. In other words, the observation function $\mathcal{O}:\mathcal{S}\times\{1,\dots,n_{a}\}\rightarrow\Omega$ provides each agent $i$ with a local observation $o^{i}$ based on the next state $s^{\prime,i}$. The joint observation can be defined as $\textbf{o}=\left\langle o^{1},\cdots,o^{n_{a}}\right\rangle$ and each agent shares the same reward function $\mathcal{R}(s,\textbf{a})\in\mathbb{R}$. Each agent has policy $\pi^{i}$ and the joint policy is denoted as $\mathbf{\pi}=\ \left\langle\pi^{1},\cdots,\ \pi^{n_{a}}\right\rangle$. The optimal joint policy is found through maximizing the joint expected reward $\mathbb{E}\left[\sum_{t=0}^{t=T}\gamma^{t}r_{t}\right]$ where $r_{t}\ =\ \mathcal{R}\left(s_{t},\ a_{t}\right)\ \in\mathbb{R}$ at each time step $t$ where $T$ is the time horizon and the discount factor $\gamma\in\left[0,1\right]$.
Each inventory control agent within the system is characterized by the following attributes:
State Space. In the context of our inventory control system, the observation set for each agent $i$ is $o^{i}=[v,b,p,d^{-1},\dots,d^{-M},o_{r}^{-1},\dots,o_{r}^{-M}]$ where $o^{i}\ \in\Omega^{i}$. In the observation set, a new variable is introduced, $p$, which is the pipeline inventory equal to the sum of order replenishment that has not yet arrived at the node from other upstream nodes.
@@
To mitigate the problem of partial observability, we include demand history and order history up to $M$ time-steps in the past where $M$ is a hyperparameter. While Recurrent Neural Networks (RNNs) could handle sequential data and temporal dependencies, we chose not to use them here for simplicity and practicality in training. Our decision to include a fixed window of past observations instead allows for a simpler implementation while still capturing relevant historical context. It is important to note that this inclusion of historical data introduces a violation of the Markov property, which states that the future state of the system only depends on the current state and not on the sequence of events that preceded it. However, real-world decision-making processes rarely exhibit perfect Markovian behavior. Therefore, including historical data is a well-established approach to augment the observation space in partially observable environments (Liu et al., 2022b; Mousa et al., 2023; Uehara et al., 2022).
Action Space. The action space is traditionally modeled as the order replenishment quantity, $o_{r}$. While the actual order replenishment quantities in our environment are discrete, we model the action space as continuous in $\left[-1,1\right]$ for scalability to a wider range of possible order sizes. In this paper we parameterize a heuristic inventory policy, specifically an $(s,S)$ policy, using a neural network policy. The $(s,S)$ policy is defined by two key parameters: the reorder point $s$, which triggers a replenishment order when the inventory reaches a specific level, and the order-up-to level $S$, which is the target inventory level after replenishment. The replenishment quantity, $o_{r}$, is dynamically determined as the difference between the order-up-to level and the current inventory level. Unlike traditional implementations that rely on fixed values for $s$ and $S$, we extend this approach by parameterizing $s$ and $S$ as stochastic variables drawn from Gaussian distributions, allowing for dynamic adaptation to demand and lead-time uncertainties.
For each agent $i$, at each time step $t$, the policy outputs the mean and standard deviation for both the reorder point $s$ and order-up-to level $S$. The policy is defined as follows:
$\pi_{i}(o^{i})=(\mu_{s_{\text{inv},t}^{i}},\sigma_{s_{\text{inv},t}^{i}},\mu_{S_{\text{ord},t}^{i}},\sigma_{S_{\text{order},t}^{i}})$ (18)
where the reorder point $s_{\text{inv},t}^{i}$ and the order-up-to level $S_{\text{ord},t}^{i}$ are sampled from a Gaussian distribution:
$s_{\text{inv},t}^{i}\sim\mathcal{N}(\mu_{s_{\text{inv},t}^{i}},\sigma_{s_{\text{inv},t}^{i}}^{2})$ (19)
$S_{\text{ord},t}^{i}\sim\mathcal{N}(\mu_{S_{\text{ord},t}^{i}},\sigma_{S_{\text{ord},t}^{i}}^{2})$ (20)
A min-max post processing step is then used to scale the values to a suitable range denoted by a subscript $s$, leading to $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ where $s_{t_{\text{inv,s}}}^{i}$ is the reorder point and $S_{t_{\text{ord,s}}}^{i}$ is the order-up-to-level. The min-max scaling step is defined as:
$f(s,s_{\text{min}},s_{\text{max}})$ $=\frac{s+1}{2}\times(s_{\text{max}}-s_{\text{min}})+s_{\text{min}}$ (21)
$s_{t_{\text{inv,s}}}$ $=f(s_{\text{inv},t},s_{\text{inv, max}},s_{\text{inv, min}})$ (22)
$S_{t_{\text{ord,s}}}$ $=f(S_{\text{ord},t},S_{\text{ord, max}},S_{\text{ord, min}})$ (23)
where $s_{\text{inv,min}}$, $s_{\text{inv,max}}$, $S_{\text{ord,min}}$ and $S_{\text{ord,max}}$ represent the lower and upper bounds for $s$ and $S$ respectively.
Once scaled, the reorder point $s_{t_{\text{inv,s}}}^{i}$ and the order-up-to level $S_{t_{\text{ord,s}}}^{i}$ are used to determine the order quantity. When the inventory reaches a level of $s_{t_{\text{inv,s}}}^{i}$, an inventory order is placed where $o_{r,t}^{i}=S_{t_{\text{ord,s}}}^{i}-v_{t}^{i}$ , rounded to the nearest integer. The neural network architecture for each actor, including the post-processing step that results in the order replenishment quantity, is illustrated in Figure 4.